Sketch the phase portrait for each linear system below

Y. dt - (33) = (211) dY dt Y.
Answer #1

Part 1

classmate Date Page d? et A AS A is diagonal matrix. and eigenvalues are 2,=3 22=3 0 are and v= and = to A. eigenve ctors of

See clear graph here

4G 68% 2:16 4.5 4. UN 3.5 3 2.5 2 1.5 1 0.5 0 -0.5 -1 -1.5 E -2 -2.5 -3 -3.5 -4 -4.5 5 -4.5 -4 -3.5 -3 -2.5 -2 -1.5 -1 -0.5 0

Part 2

classmate Date Page dr dt 2 4 A- [14] A-03-D: 41-3 - 2-> 4-2 det CA-DI) = (2-7) (4-2) +1 9 -6.atae A are a=3 and 72 3 Eigenva​​​​​see graph

4G 2 4 63% 2:35 4.5 4 3.5 IN 3 2.5 2. 1.5 1 0.5 0 -0.5 -1 -1.5 -2 -25 -3 -3.5 -4 -4.5 -4.5 -3.5 -3 -2.5 -2. -1.5 -1 -0.5 0 0.